OpenStudy (anonymous):

Estimate the quotient of eight and one over five divided by one and eight over nine using compatible numbers. What is the quotient? 4 2 8 9

OpenStudy (asevilla5):

8 1/5 is close to 8 and 1 8/9 is close to 2. Divide 8 by 2 to obtain 4. Check: Let's do the actual division for comparison. 8 1/5 = (5 * 8 + 1)/5 = 41/5 1 8/9 = (9 * 1 + 8)/9 = 17/9 Hence, [8 1/5]/[1 8/9] = [41/5]/[17/9] = [41/5] * [9/17] = 369/85 = 4.34 to the nearest hundredth.

OpenStudy (asevilla5):

so it would be 4